HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ACCESSING HIGHER VIBRATIONAL STATES OF He_{2}^{+} THROUGH MULTI-STEP EXCITATION

\underline{M. HOLDENER} , F. MERKT, Department of Chemistry and Applied Biosciences, ETH Zürich, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland

The precise measurement of highly excited vibrational states of small molecular systems such as H_{2}^{+}, H_{2} and He_{2}^{+} is of significant interest as benchmark to ab-initio quantum-chemical calculations ^{1}. In the case of He_{2}^{+}, only the first few vibrational levels of the electronic ground state X^{+} ^{2}Σ_{u}^{+} were precisely measured ^{2}^{,}^{3}. In this poster presentation, we propose a method to access the higher vibrational states of He_{2}^{+} X^{+} ^{2}Σ_{u}^{+} using a multi-step excitation scheme. Our approach involves the production of a molecular beam of He_{2} in the long-lived metastable a^{3}Σ_{u}^{+} state^{4} through an electric discharge. We then utilize a high-intensity laser to promote the system to the electronic state c ^{3}Σ_{g}^{+} in an excited vibrational level with v in the range 3 − 5. This state predominantly decays radiatively to the a ^{3}Σ_{u}^{+} state with v′ = v because of favorable Frank Condon factors. A second laser is then employed to induce a transition from the vibrationally excited metastable state a ^{3}Σ_{u}^{+}(v′) to the ion state X^{+} ^{2}Σ_{u}^{+}(v^{+} = v′). This method offers the prospect of studying the structure and dynamics of highly excited vibrational levels of He_{2}^{+} that are typically challenging to access.

1doi:10.1103/PhysRevLett.125.213001, Dávid Ferenc et al., Phys. Rev. Lett., 125, 213001 (2020)

2doi:10.1103/PhysRevLett.124.213001, Luca Semeria et al., Phys. Rev. Lett., 124, 213001 (2020)

3doi:10.1063/1.5051089, Paul Jansen et al., J. Chem. Phys., 149, 154302 (2018)

4doi:10.1063/1.455993, Cary F. Chabalowski et al., J. Chem. Phys., 90, 2504-2512 (1989)

p-number: p036  Submitted on Mon, 05 Jun 2023 17:38:43 +0200